# BOOP: Write Right Code

### v1

@@ Abstract
Novice programmers frequently adopt a syntax-specific and test-case-driven approach, writing code first and adjusting until programs compile and test cases pass, rather than developing correct solutions through systematic reasoning. AI coding tools exacerbate this challenge by providing syntactically correct but conceptually flawed solutions. In this paper, we introduce BOOP (Blueprint, Operations, OCaml, Proof), a structured framework requiring four mandatory phases: formal specification, language-agnostic algorithm development, implementation, and correctness proof. This shifts focus from “making code work” to understanding why code is correct.
BOOP was implemented at our institution using a VS Code extension and preprocessor that enforces constraints and identifies counterproductive patterns. Initial evaluation shows improved algorithmic reasoning and reduced trial-and-error debugging. Students reported better edge case understanding and problem decomposition, though some initially found the format verbose. Instructors observed stronger foundational skills compared to traditional approaches.
Keywords:
Correct-by-Construction Programming Formal Methods Program Design Programming Proofs Programming Tools
@@ 1 Introduction
Computer Science, at its core, is a discipline of abstraction, logic, and problem-solving. Writing code is a powerful way to express these ideas. Yet, we frequently note that students beginning their journey in computer science at our institution conflate the tool with the discipline and adopt a predominantly program-centric mindset. This manifests in several ways: students incorrectly confuse the expected behaviour of the program on a small set of test cases with total functional correctness, attempt to mechanically patch incorrect programs instead of reasoning from first principles, and develop a syntax-specific mindset in programming. It has been observed that students even tend to overestimate their understanding of program constructions [25]. This is reflected in students’ prioritisation of coding a solution over understanding the underlying algorithm. Roundabout approaches often result in students only being able to consider test cases that are over-fitted to their code, disregarding the problem’s specifications and total functional correctness.
This is more prominent in the age of AI-based programming tools and co-pilots. Students unfamiliar with the computational thinking process are tempted to consult language models for answers to programming problems [34]. However, language models churn out poor-quality code, and sometimes even inaccurate statements, which leads to unfit programming practices [33, 3]. This recursively worsens the student’s skill in and outlook on programming. Students also develop misconceptions about the necessity of learning problem-solving skills when LLMs could generate solutions for them, and even trust LLMs much more than LLMs’ actual accuracy calls for [33]. The ACM 2023 guidelines for undergraduate computer science education explicitly recommend instruction in proof-writing and correctness verification, including loop invariant application for algorithm correctness and inductive methods for recursive algorithm analysis [20].
@@ 2 Related Work
Computing education research has extensively documented the challenges students face when learning introductory programming. Research indicates that these difficulties stem not only from the conceptual complexity of programming constructs but also from significant barriers in the practical application of programming knowledge [25]. Students step in with preconceptions about programming difficulty or develop negative attitudes early in their learning journey, leading to ineffective learning approaches such as rote memorization and, in some cases, academic dishonesty.
A common pedagogical intervention is reducing extraneous cognitive load while preserving the essential problem-solving aspects of programming instruction. Constrained environments like mini-languages (turtle graphics) and block-based systems (MIT’s LogoBlocks [4]) mimic basic programming concepts like achieving a goal by breaking it down into a correct sequence of instructions. These approaches use syntactic constraints as scaffolding and support “hybrid programming environments that blend block-based and text-based tools” [37]. Visualization tools and other adaptive systems offer personalized feedback on simplified representations of problem-solving, objects, and functions [14, 7, 18].
However, fundamental programming concepts must not be over-abstracted such that application and formalization become disconnected. While referential frameworks can provide an initial conceptual understanding, students’ learning and evaluation should be guided by a more robust and well-structured system that comprises both formal and informal methods.
@@ 2.2 AI and CS Pedagogy
Recent research has examined the intersection of artificial intelligence tools and functional programming education, revealing significant pedagogical challenges. While large language models perform well with high-resource languages like Python, their capabilities with low-resource languages such as OCaml remain limited [6]. Even when ChatGPT provides correct solutions to functional programming problems in OCaml 68% of the time, only half of these solutions are pedagogically valuable or "benefit students in some form"[33]. Pedagogy can leverage the fact that ChatGPT-generated code exhibits detectable patterns that even remain consistent across prompt variations[6]. Given students’ inclination toward adopting AI tools, researchers recommend emphasizing code comprehension, debugging, and correctness within structured design frameworks [34, 10].
@@ 2.3 Pattern-Based Approaches to Programming Education
Educational research has also explored pattern-based pedagogical frameworks such as POSA and GoF [5, 11]. These patterns encapsulate successful solution strategies by describing problem contexts, competing forces, and balancing structures. Owen Astrachan of Duke University recommends implementing pattern-based instruction in introductory computer science courses [2] to promote "reusable thinking", encouraging students to rely on established structure over intuition. Although these ideas are valuable, they take away from the autonomy and worthwhile struggle of a student. A pedagogical approach must emphasize independent pattern discovery and understanding over pattern memorization, prioritizing the development of transferable algorithmic thinking skills.
@@ 2.4 Correctness-by-Construction
Several frameworks for stable, correct-by-construction programming have emerged. Meyer’s Design By Contract (DbC) is a systematic approach to designing programs where preconditions, postconditions, and invariants are explicitly defined for each component [29]. Wing explicitly mentions that computational thinking involves using “invariants to describe a system’s behaviour succinctly and declaratively” [38]. Gries and Schneider emphasize the use of a rigorous, equational style of proving and show how logic can be applied to formally develop programs from specifications [13, 12]. Loksa et al. investigated meta-cognitive awareness in novice programmers by identifying six problem-solving stages and explicitly coaching students to recognize which stage they occupied when encountering difficulties [28]. A study which leveraged their framework to investigate common errors in novices found that successful students differed primarily in forming correct initial problem conceptualization, corresponding to the first stage: reinterpret the problem prompt [32]. There is a need for pedagogical approaches that integrate formal specification methods with novice-friendly structured learning environments.
@@ 3 Overview
In this section, we demonstrate our framework through a representative programming task from Introduction to Computer Science at our institution. The selected exercise exemplifies the cognitive challenges inherent in introducing formal specification-driven programming to novice computer science students:
@@ 3.3 Blueprint Specification
Addressing this epistemological shift requires explicit mathematical specification as a pedagogical intervention. As Lamport (2015) provocatively asks, “Who builds a house without drawing a blueprint?.” Yet programmers routinely construct complex systems without formal specifications [26, 35, 30, 31]. Research in CS education consistently demonstrates that students who engage with formal preconditions and postconditions develop robust problem-solving skills [29, 38, 13].
@@ 3.4 Algorithm as Operations
This solution handles all cases appropriately, maintains readability, and lends itself to formal verification. One can easily adapt these operations for imperative implementations using loops and mutable variables if functional programming is not the target paradigm (as shown in Appendix 0.C). Moreover, this algorithmic structure generalizes beyond natural numbers to any Euclidean domain where division with remainder is well-defined. The core operations of comparison, subtraction, and iterative reduction remain conceptually identical.
@@ 3.5 Proof of Correctness
This theorem asserts that the program satisfies its specification (blueprint) in all well-defined cases. To establish this, we proceed by structural induction on a : nat, analysing the recursive structure of the function and verifying that both postconditions are preserved at each step. A complete proof, as expected of students in our Introduction to Computer Science course, is provided in Appendix 0.B.
@@ 4 The BOOP Framework for Systematic Problem-Solving
This shift toward human-centric program development aligns with our pedagogical goals of fostering language-agnostic algorithmic thinking. Literate programming principles emphasize structured design with clear separation of concerns, human-first ordering that prioritizes cognitive comprehension over execution sequence, and meta-language abstraction that builds logical constructs before implementation details.
@@ 4.1 The Four-Phase BOOP Framework
The BOOP acronym, where the second O stands for ‘OCaml code’ was chosen because the Introduction to Computer Science course at our institution is taught in OCaml, and because of the acronym’s appeal and memorability among students. Currently, BOOP supports OCaml in its initial deployment; however, it is a language-agnostic framework for problem-solving, as demonstrated with an example of Java code in Appendix 0.C.
@@ 4.2.1 Program Preprocessor
Based on experience from Spring 2025, students increasingly submit AI-generated code that circumvents learning objectives while appearing syntactically correct. Our initial attempts at surveillance mechanisms proved largely ineffective. The experience motivated us to develop a program preprocessor that promotes conceptual understanding through instructive feedback. This tool parses OCaml submissions before compilation, identifying counterproductive patterns (such as the use of built-in functions like List.rev instead of native implementations, imperative constructs like mutation and variable reassignment that contradict functional programming paradigms, nested functions defined in local scope that obscure algorithmic structure, and functions with anonymous inputs that prevent clear reasoning about program behaviour.
Furthermore, our system restricts specific syntax and practices that undermine learning goals: it encourages recursive thinking over imperative loops, mandates explicit type reasoning through typed function definitions, promotes first-principles implementation over excessive built-in function usage, and requires pattern matching where it would be more appropriate than if-then-else control flow. These constraints are configurable and can be dynamically adjusted based on assignment objectives, student proficiency levels, and classroom instruction progress. As new concepts are introduced in lectures, the preprocessor is updated, creating a responsive system that evolves with the curriculum.
@@ 4.2.2 Integrated Development Environment
The BOOP framework and preprocessor are integrated through a custom VS Code extension that provides comprehensive support. The extension features an instructor interface for defining assignment parameters and allowed syntactic patterns, specialized syntax highlighting for BOOP section delineation, a check ensuring all four phases are present, and structured evaluation tools that enable targeted feedback on conceptual understanding rather than mere input-output correctness. Snapshots of the IDE and a link to download the bundle are available at Appendix 0.D).
This integrated environment transforms the traditional code-compile-test cycle into a specification-design-implement-verify workflow. By embedding pedagogical constraints directly into the development environment, we create a learning ecosystem where good practices become the path of least resistance, fostering sustainable computational thinking skills.
@@ 5.2 Implementation Challenges and Student Misconceptions
As a part of the evaluation, we consulted CS educators with decades of experience in teaching introductory courses. All of the consulted educators recognised value in our framework as it encouraged rigorous articulation of requirements and set the appropriate tone and momentum for a computer science trajectory. Instructors in advanced programming courses offered in our institution (and otherwise [8]) encounter students with inadequate programming practices and weak methods. Early intervention through approaches like BOOP may address these persistent educational challenges and make students well-prepared for future coursework.
@@ 6 Conclusion and Future Directions
    Integrate BOOP with automated theorem proving and verification environments capable of generating counterexamples and evaluating student-constructed proofs. Tools such as Dafny, Lean, or Why3 could provide immediate feedback on proof correctness while automatically generating edge cases that challenge student solutions, to establish a comprehensive pipeline supporting automated assessment and systematic feedback.
@@
    Augment BOOP with an intelligent tutoring system that can provide individualized scaffolding and complement the supportive role of teaching assistants. This work is motivated by a study at Harvard that demonstrated substantial success with LLM-based tutoring systems, reporting high accuracy in responses and positive student perceptions of personalized instruction [27].
  3. 3.
    Extend BOOP to incorporate additional analytical components that deepen students’ understanding of asymptotic computational complexity as well as unit testing. We will also add support for additional programming languages commonly used in computer science education, including Java and C++.
  4. 4.
    We have planned for a comprehensive longitudinal study planned for the Monsoon 2025 semester. We continue to explore multi-institutional deployments, comparative analysis with traditional programming pedagogies, longer-term retention studies, and industry preparation assessment.
@@ Appendix 0.B Proof of Correctness for the Division Function
      $a - b = plus (mult b q') r'|,\quad\text{and}\quad\verb|r' < b|.$
@@
Appendix 0.C An Imperative Division Algorithm
This appendix demonstrates how the operational steps developed in section˜3.4 can be translated into an imperative paradigm. For clarity and convenience, we will use the arithmetic operators +, *, and < with the nat operations as presented in Listing˜1.
The blueprint remains the same:
To write a correct div, the function contract is:
  1. 1.
    Requires: a : nat, b : nat, and b <> Zero (that is b is non-zero).
  2. 2.
    Ensures: Returns a pair (q, r) such that:
    1. (a)
      a = (b * q) + r
    2. (b)
      less_than r b = true
The imperative procedure is derived by interpreting the first postcondition, a = (b * q) + r, as an invariant. This invariant holds with the initial assignments q := Zero and r := a. Since division can be understood as repeated subtraction, the procedure repeatedly subtracts b from r until the second postcondition, r < b, is satisfied. The steps are as follows:
  1. 1.
    Check if b = Zero. In this case return a runtime error.
  2. 2.
    If not, proceed by setting variables q set to Zero and r set to a.
  3. 3.
    Until we have r < b, repeatedly subtract b from r and each time update q := Succ !q.
  4. 4.
    When r < b, return the values for q and r.
The operations can be translated to code. To emphasise on the language-agnostic nature of our approach, we provide the code in OCaml syntax in Listing˜3 and Java syntax in Listing˜4.
1let minus (n : nat) (m : nat) : nat =
2 let n_ref = ref n in
3 let m_ref = ref m in
4
5 while !m_ref <> Zero do
6 if !n_ref = Zero then failwith "minus: function requires n >= m"
7 else
8 match !n_ref, !m_ref with
9 | Succ n’, Succ m -> n_ref := n’; m_ref := m
10 | _ -> assert false
11 done;
12 !n_ref
13
14let div (a : nat) (b : nat) : (int * int) =
15 match b with
16 | Zero -> failwith "Division by zero"
17 | _ ->
18 let q = ref Zero in
19 let r = ref a in
20 while not (less_than !r b) do
21 r := minus !r b;
22 q := Succ !q
23 done;
24 (!q, !r)
Listing 3: A translation of the operational steps for imperative division and safe subtraction on natural numbers in OCaml syntax. We assume the availability of a correct less_than function for brevity.
1public class Peano {
2
3 abstract static class Nat {
4 abstract boolean isZero();
5 }
6
7 static final class Zero extends Nat {
8 @Override boolean isZero() { return true; }
9 }
10
11 static final class Succ extends Nat {
12 final Nat prev;
13 Succ(Nat prev) { this.prev = prev; }
14 @Override boolean isZero() { return false; }
15 }
16
17 static Nat minus(Nat n, Nat m) {
18 while (!m.isZero()) {
19 if (n.isZero()) throw new IllegalArgumentException("minus: n < m");
20 n = ((Succ) n).prev;
21 m = ((Succ) m).prev;
22 }
23 return n;
24 }
25
26 static Nat[] div(Nat a, Nat b) {
27 if (b.isZero()) throw new ArithmeticException("division by zero");
28 Nat q = new Zero();
29 Nat r = a;
30 while (!lessThan(r, b)) {
31 r = minus(r, b);
32 q = new Succ(q);
33 }
34 return new Nat[] { q, r };
35 }
36}
Listing 4: A translation of the operational steps for division and safe subtraction on natural numbers in Java syntax. We assume the availability of a correct lessThan function for brevity.
The correctness of the implementation can be presented as a theorem:
Theorem 0.C.1
  1. 1.
  2. 2.
    If b <> Zero, then div a b returns a pair (q, r) such that:
    1. (a)
      a = (b * q) + r, and
    2. (b)
      r < b.
Proof
We analyze the correctness of the implementation using loop invariants.
Case 1: b = Zero.
The implementation immediately raises an exception:
match b with Zero -> failwith "Division by zero"
Thus, the behavior matches the specification.
Case 2: b <> Zero.
The function proceeds with the following imperative structure:
let q = ref Zero in let r = ref a in while not (less_than !r b) do   r := minus !r b;   q := Succ !q done; (!q, !r)
We define a loop invariant that must hold at the beginning of each iteration of the loop: a = (b * !q) + !r  and  !r : nat.
Initialization: Before entering the loop, we have q = Zero and r = a. Then,
(b * !q) + !r = (b * Zero) + a = Zero + a = a
Hence, the invariant holds before the first iteration.
Maintenance: Assume the invariant holds at the start of an iteration. Then, inside the loop, we update r := minus !r b and q := Succ !q. Let q_old = !q before the iteration and r_old = !r before the iteration. Then after the iteration we have:
q := Succ q_old r := minus r_old  b
Using the assumption that r < b = false, the subtraction is valid, and:
a = (b * q_old) + r_old = (b * (Succ q_old)) + (r_old - b) = (b * q!) + r!
So the invariant continues to hold after each iteration.
Termination: The loop terminates when !r < b. At this point, the invariant still holds. Hence the returned pair (!q, !r) satisfies the postconditions. ∎
Appendix 0.D Preview of the IDE
Figure 1: Screenshots of the VS Code IDE with the ‘BOOP: How to Right Code’ (version 0.5.9) Extension enabled
The complete bundle can be found at zenodo.org/records/16487904. The current version (0.5.9) is a preliminary one, and version 1.0 is under development.

### v2

@@ Abstract
Novice programmers frequently adopt a syntax-specific and test-case-driven approach, writing code first and adjusting until programs compile and test cases pass, rather than developing correct solutions through systematic reasoning. AI coding tools exacerbate this challenge by providing syntactically correct but conceptually flawed solutions. In this paper, we address the question of developing correctness-first methodologies to enhance computational thinking in introductory programming education. To this end, we introduce BOOP (Blueprint, Operations, OCaml, Proof), a structured framework requiring four mandatory phases: formal specification, language-agnostic algorithm development, implementation, and correctness proof. This shifts focus from making the code work to understanding why the code is correct.
BOOP was implemented at Ashoka University using a VS Code extension and preprocessor that enforces constraints. Initial evaluation shows improved algorithmic reasoning and reduced trial-and-error debugging. Students reported better edge case understanding and problem decomposition, though some initially found the format verbose. Instructors observed stronger foundational skills compared to traditional approaches, suggesting that structured correctness-first approaches may significantly improve students’ computational thinking abilities.
@@ 1 Introduction
Computer Science, at its core, is a discipline of abstraction, logic, and problem-solving. Writing code is a powerful tool to express these ideas. Yet, we frequently note that students beginning their journey in computer science at Ashoka University conflate the tool with the discipline and adopt a predominantly program-centric mindset. This manifests in several ways: students incorrectly confuse the expected behaviour of the program on a small set of test cases with total functional correctness, attempt to mechanically patch incorrect programs instead of reasoning from first principles, and develop a syntax-specific mindset in programming. In general, students even tend to overestimate their understanding of program constructions [25]. This is reflected in their prioritisation of coding a solution over understanding the underlying algorithm. Roundabout approaches often result in students only being able to consider test cases that are over-fitted to their code, disregarding the problem’s specifications and overall correctness. The ACM 2023 guidelines for undergraduate computer science education explicitly recommend instruction in proof-writing and correctness verification, including loop invariant application for algorithm correctness and inductive methods for recursive algorithm analysis [24].
This is more prominent in the age of AI-based programming tools and co-pilots. Students unfamiliar with the computational thinking process are tempted to consult language models for answers to programming problems [34]. However, language models churn out poor-quality code, and sometimes even inaccurate statements, which leads to poor programming practices [32, 3]. This worsens the student’s skill in and outlook on programming. They develop misconceptions about the necessity of learning problem-solving skills when tools such as Large Language Models (LLMs) could generate solutions for them, trusting these tools much more than they call for [32].
@@ 2 Related Work
Computing education research has extensively documented the challenges students face when learning introductory programming. Students step in with preconceptions about programming difficulty or develop negative attitudes early in their learning journey, leading to ineffective learning approaches such as rote memorization and, in some cases, academic dishonesty.
A common pedagogical intervention is reducing extraneous cognitive load while preserving the essential problem-solving aspects of programming instruction. Constrained environments like mini-languages (turtle graphics) and block-based systems (MIT’s LogoBlocks [4]) mimic basic programming concepts like achieving a goal by breaking it down into a correct sequence of instructions. These approaches use syntactic constraints as scaffolding and support “hybrid programming environments that blend block-based and text-based tools” [37].
However, fundamental programming concepts must not be over-abstracted to the point that application and formalization become disconnected. While referential frameworks can provide an initial conceptual understanding, students’ learning and evaluation should be guided by a more robust and well-structured system that comprises both formal and informal methods.
@@ 2.2 AI and CS Pedagogy
Recent research has examined the intersection of artificial intelligence tools and programming education, revealing significant pedagogical challenges. While LLMs perform well with high-resource languages like Python, their capabilities with low-resource languages such as OCaml remain limited [6]. Even when ChatGPT provides correct solutions to functional programming problems in OCaml 68% of the time, only half of these solutions are pedagogically valuable or “benefit students in some form”[32]. Moreover, ChatGPT-generated code exhibits detectable patterns that even remain consistent across prompt variations[6]. Accordingly, researchers recommend emphasizing code comprehension, debugging, and correctness within structured design frameworks [34, 10].
@@ 2.3 Pattern-Based Approaches to Programming Education
Educational research has also explored pattern-based pedagogical frameworks such as POSA and GoF [5, 11]. These patterns encapsulate successful solution strategies by describing problem contexts, competing forces, and balancing structures. Although these ideas are valuable, they take away from the autonomy and worthwhile struggle of a student. A pedagogical approach must emphasize independent pattern discovery and understanding over pattern memorization, prioritizing the development of transferable algorithmic thinking skills.
@@ 2.4 Correctness-by-Construction
Several frameworks for stable, correct-by-construction programming have emerged. Meyer’s Design By Contract is a systematic approach to designing programs where preconditions, postconditions, and invariants are explicitly defined for each component [29]. Wing explicitly mentions that computational thinking involves using “invariants to describe a system’s behaviour succinctly and declaratively” [38]. Gries and Schneider emphasize the use of a rigorous, equational style of proving and show how logic can be applied to formally develop programs from specifications [14, 13]. Loksa et al. investigated meta-cognitive awareness in novice programmers by identifying six problem-solving stages and explicitly coaching students to recognize which stage they occupied when encountering difficulties [28]. A study which leveraged their framework to investigate common errors found that successful students differed primarily in forming correct initial problem conceptualization, corresponding to the first stage: reinterpret the problem prompt [33]. There remains a need to integrate formal specification methods with novice-friendly structured learning environments for CS education.
@@ 3 Overview
In this section, we demonstrate our framework through a representative programming task from the Introduction to Computer Science course at our institution. The selected exercise exemplifies the challenges inherent in introducing formal specification-driven programming to CS students:
@@ 3.3 Blueprint Specification
Addressing this shift requires explicit mathematical specification as a pedagogical intervention. Leslie Lamport compares the lack of specifications to building “a skyscraper without a blueprint” [26]. And yet, programmers routinely construct complex systems without formal specifications [26, 35, 30, 31]. Research in CS education consistently shows that students who engage with formal preconditions and postconditions develop robust problem-solving skills [29, 38, 14].
@@ 3.4 Algorithm as Operations
This solution handles all cases appropriately, maintains readability, and lends itself to formal verification. One can easily adapt these operations for imperative implementations using loops and mutable variables if functional programming is not the target paradigm. Moreover, this algorithmic structure generalizes beyond natural numbers to any Euclidean domain where division with remainder is well-defined. The core operations of comparison, subtraction, and iterative reduction remain conceptually identical.
@@ 3.5 Proof of Correctness
This theorem asserts that the program satisfies its specification (blueprint) in all well-defined cases. To establish this, we proceed by structural induction, analysing the recursive structure of the function and verifying that both postconditions are preserved at each step. A complete proof, as expected of students in our Introduction to Computer Science course, is provided in Appendix 0.B.
@@ 4 The BOOP Framework for Systematic Problem-Solving
Literate programming emphasizes structured design with clear separation of concerns, human-first ordering that prioritizes cognitive comprehension over execution sequence, and meta-language abstraction that builds logical constructs before implementation details. This shift toward human-centric program development aligns with our goals of fostering language-agnostic algorithmic thinking.
@@ 4.1 The Four-Phase BOOP Framework
The BOOP acronym, where the second O stands for ‘OCaml code’ was chosen because the course at our institution is taught in OCaml, and because of the acronym’s appeal and memorability among students. Currently, BOOP supports OCaml but it is a language-agnostic framework for problem-solving and can be adapted to support other languages.
@@ 4.2.1 Program Preprocessor
Based on experience from Spring 2025, students have been increasingly submitting AI-generated code. This circumvents learning objectives while appearing syntactically correct. The experience motivated us to develop a preprocessor that promotes conceptual understanding through instructive feedback. Our tool parses OCaml submissions before compilation, identifying counterproductive patterns (such as the use of built-in functions like List.rev instead of native implementations, imperative constructs like mutation and variable reassignment that contradict functional programming paradigms, nested functions defined in local scope that obscure algorithmic structure, and functions with anonymous inputs) that prevent clear reasoning about program behaviour.
Furthermore, our system restricts specific syntax and practices that undermine learning goals: it encourages recursive thinking, mandates explicit type reasoning through typed function definitions, promotes first-principles implementation over excessive built-in function usage, and requires pattern matching where it would be more appropriate than if-then-else control flow. These constraints are configurable and can be dynamically adjusted based on assignment objectives, student proficiency levels, and classroom instruction progress. However, while the framework allows for substantial customisation by the instructor, it is not a requirement. Assignments can be deployed with liberal default settings.
@@ 4.2.2 Integrated Development Environment
On the technical side, the only requirement is installing VS Code, which we found was already the preferred IDE for most students. A custom VS Code extension integrates the BOOP framework with an instructor interface for configuring assignments, specialized syntax highlighting, a completion check to enforce all four phases, and evaluation tools for targeted, conceptual feedback.
The integrated environment transforms the traditional code-compile-test cycle into a specification-design-implement-verify workflow. We create a learning ecosystem where good practices become the path of least resistance, fostering sustainable computational thinking skills.
@@ 5.2 Implementation Challenges and Student Misconceptions
As a part of the evaluation, we also consulted CS educators with decades of experience in teaching introductory courses. All of the consulted educators recognised value in our framework as it encouraged rigorous articulation of requirements and set the appropriate tone and momentum for a computer science trajectory. Instructors in advanced programming courses offered in our institution (and otherwise [8]) encounter students with inadequate programming practices and weak methods. Early intervention through approaches like BOOP may address these persistent educational challenges and make students well-prepared for future coursework. We continue to explore multi-institutional deployments, comparative analysis with traditional programming pedagogies, longer-term retention studies, and industry preparation assessment.
@@ 6 Conclusion and Future Directions
    Integrate BOOP with automated theorem proving and verification environments capable of generating counterexamples and evaluating student-constructed proofs. Tools such as Dafny, Lean, or Why3 are scalable and provide immediate feedback on proof correctness while automatically generating edge cases that challenge student solutions, to establish a comprehensive pipeline supporting automated assessment and systematic feedback. As an easier alternative, be combined with online judges for test-case verification, coming closer to current traditional modes of checking the accuracy of code against a problem statement.
@@
    To facilitate framework adoption after an initial phase of resistance among students, we plan to tailor lectures to mirror the problem-solving paradigm of BOOP, in an attempt to discourage the notion that BOOP is a final formatting step for assignment submissions only. The perception of BOOP as verbose often stems from students who reverse-engineer correctness criteria, leading to repetition rather than genuine problem analysis. To address this, lectures must emphasise that BOOP components are problem-solving tools, not mere submission requirements. Since students have favored instant feedback, a potential direction is also to use non-suggestive, LLM-driven prompts. We can rate-limit these to discourage frivolous use.
@@
Acknowledgements.
We gratefully acknowledge the Mphasis AI & Applied Tech Lab at Ashoka University for supporting this research.
@@ Appendix 0.B Proof of Correctness for the Division Function
[…]